OpenStudy (anonymous):

Given the parent functions f(x) = 3x + 2 and g(x) = 5x − 10, what is g(x) − f(x)?

OpenStudy (tkhunny):

Have you considered subtraction?

OpenStudy (anonymous):

don't forget parentheses what the monkey is a "parent function" ??

OpenStudy (solomonzelman):

Consider that f(x) = 3x + 2 g(x) = 5x − 10 OK?

OpenStudy (anonymous):

Ok go on i'm getting this.

OpenStudy (solomonzelman):

f(x) - g(x) just means subtract what g(x) is equal to from what f(x) is equal to.

OpenStudy (solomonzelman):

rewrite f(x)-g(x) in numbers for me really quick please.

OpenStudy (anonymous):

so (3x+2)-(5x-10)

OpenStudy (solomonzelman):

THERE YOU GO!

OpenStudy (anonymous):

Ok! yes let me continue I think you hit my smart string

OpenStudy (solomonzelman):

Ok, just in case, when you are done tell me what you get as the final answer.

OpenStudy (anonymous):

ok I will

OpenStudy (anonymous):

I got −2(x−6)

OpenStudy (anonymous):

@SolomonZelman

OpenStudy (solomonzelman):

Yes!

OpenStudy (anonymous):

ok so now what do I do?

OpenStudy (solomonzelman):

\(\Huge{\color{purple}{\bigstar}\color{blue}{\bigstar}\color{green}{\bigstar}\color{lime}{\bigstar}\color{cyan}{\bigstar}\color{purple}{\bigstar}\color{blue}{\bigstar}\color{green}{\bigstar}\color{lime}{\bigstar}\color{cyan}{\bigstar}\color{purple}{\bigstar}\color{blue}{\bigstar}\color{green}{\bigstar}\color{lime}{\bigstar}}\\\color{magenta }{.}\\\Huge\sf\color{blue}{GOOD!~\ddot\smile}\\\color{white}{.}\\\\\Huge{\color{purple}{\bigstar}\color{blue}{\bigstar}\color{green}{\bigstar}\color{yellow}{\bigstar}\color{cyan}{\bigstar}\color{purple}{\bigstar}\color{blue}{\bigstar}\color{green}{\bigstar}\color{yellow}{\bigstar}\color{cyan}{\bigstar}\color{purple}{\bigstar}\color{blue}{\bigstar}\color{green}{\bigstar}\color{yellow}{\bigstar}}\)

OpenStudy (solomonzelman):

You got the answer.

OpenStudy (anonymous):

wait what which one is the answer? g(x) − f(x) = 5x − 8 − 3x g(x) − f(x) = 5x − 12 − 3x g(x) − f(x) = 3x − 8 − 5x g(x) − f(x) = 3x + 12 − 5x

OpenStudy (anonymous):

@SolomonZelman

OpenStudy (solomonzelman):

(3x+2)-(5x-10) 3x+2-5x+10 -2x+12 D is most accurate, although for real, they are legit messed up, all of them.....